# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ 3.3 Consistency Transferability Estimator (CTE)
Considering a single target dataset , we rank the transfer performance of a set of models on by measuring pixelwise prediction consistency between perturbed and unperturbed predictions (Fig. 1). For classification or semantic segmentation, we can directly compare class labels of the pixels.
Firstly, we propose to extend the Effective Invariance (EI) [15], originally formulated for classification tasks, to semantic segmentation. EI measures the invariance of a single prediction to test-time perturbations, which for segmentation can be formulated as:
@@
where and represent the output of a model after softmax and its thresholded value, with representing the perturbed outputs. We refer to the CTE-EI as a “soft” metric as it measures both the consistency and confidence of the probabilistic output. This introduces a reliance on good model calibration which is not guaranteed in transfer [50]. We therefore propose another consistency measure based on the thresholded model output (“hard” consistency). This measure directly compares the number of pixelwise changes between the perturbed and unperturbed predictions, using the Hamming distance normalised by the number of pixels:
This is equivalent to the between the predictions, but considering in terms of pixel flips adds intuition. To counteract the high class imbalance of biomedical data and to allow for multiclass problems, we calculate per class consistency and then apply the weighted average. This is achieved by limiting the pixelwise iterator to the pairwise union of perturbed and unperturbed prediction masks for this class.
For instance segmentation, direct pixel comparison is impossible as instance labels are arbitrary. We therefore modify our approach to measure the consistency of assignment of pixel pairs to the same or different objects, as measured by the Rand Index [56, 67]. To account for class imbalance, we utilise the foreground-restricted Adapted Rand Score (ARS), defined by [4] as:
where is the probability that a pixel has the same label in the perturbed prediction and in the unperturbed prediction, is the probability that a pixel has label in , and is the probability that a pixel has label in .
The consistency score for a single transfer is then calculated as:
@@
where is calculated using one of the above metrics (CTE-EI, NHD, ARS). The median is used for robustness with small dataset sizes. The final ranking within is the descending order of .
@@ 3.4 Model Perturbation
Perturbations can be applied either to the model input through test-time augmentations or directly within the feature space, assuming access to intermediate model representations. Both approaches are motivated by prior work in uncertainty estimation [43, 39] and consistency training [48].
We assess prediction robustness to perturbations under transfer to OOD data, thereby probing the suitability of the embedding space and decision boundaries for target embeddings. This differs from evaluating the model’s learned invariance to perturbations within the source domain (e.g., due to data augmentation during training). Even for models trained with augmentations, perturbations can still expose informative variations on target data. To examine this, we rank models trained both with and without augmentation and consider perturbations in both input and feature spaces.
Perturbations must remain tolerable [43]: if they are too strong, even highly transferable models become inconsistent, while overly weak perturbations leave all models perfectly stable. Empirically, we find that a broad range of perturbation magnitudes satisfy this condition allowing effective transfer estimation via CTE (see Supp.).
Input Space Perturbations. As input perturbations, we explored several popular image transformations: additive Gaussian noise, gamma correction, and changes in brightness and contrast. In our experiments we found similar performance across all perturbation types. For brevity we report Gaussian noise, where controls the strength of the perturbation, except for 3D segmentation where we also report gamma results. See Supp. and Tab. 2 for additional experiments.
Feature Perturbation.
Given access to a model, perturbations can be applied directly to intermediate feature layers, while keeping the network weights frozen (greybox model). Motivated by findings in semi-supervised learning [48] and also by work in uncertainty estimation [43], we investigate test-time spatial dropout [66]. Specifically, we apply dropout (1) across all layers of the network, (2) only at the bottleneck, where information is the most redundant [25] (Tabs. 1 and 4) or (3) at the bottleneck and in the skip connection layers (Tab. 2). The dropout rate is controlled by . For a network with layers and dropout perturbation applied at layer the prediction is defined as:
@@ 4.1 Datasets
We evaluate our approach on a range of public segmentation datasets, covering six distinct tasks and multiple modalities. For each transfer prediction experiment, we fix one target dataset and evaluate the transfers from all task-specific models (see Supp. for further details).
@@
Electron Microscopy (EM), Mitochondria. We use four semantic segmentation datasets with neural tissues from different species and EM modalities: [40, 17, 54]. We also reformulated these datasets as a per-patch binary classification task, labelling patches as positive if mitochondria are present. This auxiliary task allows us to disentangle domain shift effects from the intrinsic suitability of classification-based ranking methods for semantic segmentation.
Light Microscopy (LM), Nuclei. We use nine fluorescent nuclei datasets for instance and semantic segmentation taken from a variety of species and LM modalities: [68, 69, 31, 23, 38, 11, 5, 14, 18].
@@
Light Microscopy, Cells. We consider four dense instance cell segmentation datasets, taken from fruit fly, Arabidopsis thaliana and human samples: [18, 77, 76, 53].
@@
Cone-Beam Computed Tomography (CBCT), ToothFairy2. We use the ToothFairy2 datasets taken from a MICCAI2024 challenge [9, 10], comprising 3D CBCT volumes of the human jaw with 42 labelled classes, training the models on the official challenge training set and assessing transferability on the test set acquired from another institution.
4.2 Models
@@
We trained a diverse set of model architectures for transfer comparison. For classification: ResNet18 [24], ResNet50 [24], DenseNet121 [26], MobileNetV2 [60], MobileNetV3 [55]. For binary semantic and instance segmentation: 2D U-Net [58], Residual U-Net [34] and Unetr [22]. We also ranked publicly available models such as [3], Cellpose-SAM [51], and a model from the BMZ [49], id: 5092850 ‘powerful-chipmunk‘ [53]. For multi-class semantic segmentation, we selected top-models from the ToothFairy2 challenge [72] (Isensee et al. and Wang et al.) and further included architectures covering CNNs (2D nnU-Net, residual-encoder 3D nnU-Net [28]), Transformers (SwinUNETR [21]) and State Space Models (UMamba [41], VMamba [37]).
| Task | Unsupervised | Supervised | |||||||||||||||||||
TS | CCFV | LEEP | GBC | Hscore | LEEP | LogME | NCTI | RegHscore | |||||||||
| Mito | K | 0.77 ±0.1 | 0.71 ±0.1 | 0.73 ±0.1 | 0.69 ±0.1 | 0.25 ±0.2 | -0.11 ±0.1 | 0.41 ±0.4 | 0.39 ±0.1 | 0.23 ±0.2 | 0.89 ±0.1 | 0.08 ±0.2 | 0.21 ±0.2 | 0.24 ±0.2 | |||||||
| S | 0.85 ±0.1 | 0.83 ±0.1 | 0.86 ±0.1 | 0.85 ±0.1 | 0.30 ±0.3 | -0.16 ±0.1 | 0.49 ±0.5 | 0.52 ±0.1 | 0.33 ±0.4 | 0.96 ±0.0 | 0.10 ±0.3 | 0.28 ±0.2 | 0.34 ±0.4 | ||||||||
| P | 0.83 ±0.1 | 0.79 ±0.1 | 0.86 ±0.1 | 0.84 ±0.1 | 0.23 ±0.3 | -0.18 ±0.1 | 0.39 ±0.3 | 0.47 ±0.2 | 0.37 ±0.3 | 0.97 ±0.0 | 0.06 ±0.3 | 0.37 ±0.2 | 0.37 ±0.2 | ||||||||
| Nuclei | K | 0.62 ±0.1 | 0.69 ±0.1 | 0.74 ±0.2 | 0.71 ±0.2 | 0.02 ±0.3 | -0.03 ±0.3 | 0.11 ±0.1 | -0.13 ±0.3 | 0.29 ±0.5 | 0.67 ±0.4 | -0.20 ±0.3 | 0.29 ±0.5 | 0.35 ±0.4 | |||||||
| S | 0.77 ±0.1 | 0.82 ±0.1 | 0.85 ±0.1 | 0.84 ±0.1 | 0.10 ±0.4 | -0.05 ±0.4 | 0.12 ±0.1 | -0.16 ±0.4 | 0.25 ±0.6 | 0.77 ±0.3 | -0.28 ±0.3 | 0.25 ±0.6 | 0.36 ±0.4 | ||||||||
| P | 0.98 ±0.0 | 0.97 ±0.0 | 0.90 ±0.1 | 0.62 ±0.4 | -0.09 ±0.4 | -0.16 ±0.5 | 0.32 ±0.2 | -0.08 ±0.4 | 0.31 ±0.4 | 0.94 ±0.0 | -0.49 ±0.3 | 0.30 ±0.3 | 0.34 ±0.3 | ||||||||
4.3 Baselines
We include supervised transfer metrics: LEEP [36], GBC [52], HScore [7], LEEP [47], LogME [80], NCTI [74], Regularised-HScore [27], CC-FV [79], which represent the state-of-the-art for biomedical segmentation. We also use the SOTA for unsupervised validation Transfer Score (TS) [78]. No baselines are available for instance segmentation.
4.4 Performance Evaluation
@@
For all transferability metrics, we calculate the correlation between the metric ranking and the performance score ranking. As is standard [1] we evaluate the correlation scores using: Kendall tau [30], Pearson correlation coefficient [8], and Spearman’s rank correlation coefficient [63]. The metrics vary in the range with indicating perfect positive correlation. The metrics are complementary, indicating if the agreement is influenced more by overall rank differences or local pairwise inconsistencies. While a perfect should be monotonically related to the , linear correlation is not required for transferability metrics.
@@ 5.1 Semantic Segmentation
Previous work [1] has suggested that supervised transferability metrics developed for classification tasks can be readily extended to semantic segmentation by treating it as pixelwise classification. The standard approach samples a class-balanced set of pixels from each image and applies the transfer metric directly to this per-pixel classification set. However, for our biomedical datasets, this assumption does not hold: existing transfer metrics, even with supervision, fail to predict either direct transfer performance (Tab. 1) or post–fine-tuning performance (see Supp.). To determine if this limitation arises from the biomedical imaging domain [13] or the shift to semantic segmentation, we additionally evaluated these metrics on mitochondria classification.
Fig. 2 shows the performance score (F1) plotted against each transfer score for a single target dataset (Hmito) for both classification and semantic segmentation tasks. The metrics perform well for classification (violet crosses) but, except for LEEP and our CTE, fail to reliably correlate with semantic segmentation rankings (green circles).
| Class | CTE-EI (Gauss) | CTE-NHD (Gauss) | CTE-EI (DropOut) | CTE-NHD (DropOut) | CTE-EI (Gamma) | CTE-NHD (Gamma) | |||||||
| IACs | K | 0.90 | (**) | 0.81 | (**) | 0.87 | (*) | 0.87 | (*) | 1.00 | (**) | 0.90 | (*) |
| S | 0.96 | (**) | 0.93 | (*) | 0.94 | (**) | 0.94 | (**) | 1.00 | (**) | 0.96 | (**) | |
| P | 0.99 | (**) | 0.98 | (**) | 0.83 | (*) | 0.80 | (0.1) | 0.99 | (**) | 0.98 | (**) | |
| Teeth | K | 0.81 | (**) | 1.00 | (**) | 0.67 | (0.1) | 0.75 | (0.1) | 0.71 | (*) | 0.90 | (**) |
| S | 0.89 | (*) | 1.00 | (**) | 0.77 | (0.1) | 0.77 | (0.1) | 0.86 | (*) | 0.96 | (**) | |
| P | 0.95 | (**) | 0.98 | (**) | 0.85 | (**) | 0.94 | (**) | 0.89 | (**) | 0.95 | (**) | |
| Mand. | K | 0.62 | (*) | 0.52 | (0.2) | 0.87 | (*) | 0.87 | (**) | 0.43 | (0.3) | 0.81 | (**) |
| S | 0.82 | (*) | 0.68 | (0.1) | 0.87 | (*) | 0.94 | (*) | 0.54 | (0.2) | 0.89 | (**) | |
| P | 0.73 | (0.1) | 0.65 | (0.1) | 0.89 | (**) | 0.96 | (**) | 0.74 | (0.1) | 0.80 | (*) | |
| Sinus. | K | 0.43 | (0.3) | 0.81 | (*) | 0.87 | (*) | 0.83 | (*) | 0.43 | (0.2) | 0.90 | (**) |
| S | 0.50 | (0.3) | 0.89 | (*) | 0.84 | (*) | 0.94 | (*) | 0.57 | (0.2) | 0.96 | (**) | |
| P | 0.38 | (0.4) | 0.73 | (0.1) | 0.88 | (**) | 0.90 | (**) | 0.38 | (0.4) | 0.98 | (**) | |
| Overall | K | 0.71 | (*) | 0.90 | (**) | 0.93 | (*) | 0.73 | (0.1) | 0.81 | (**) | 1.00 | (**) |
| S | 0.86 | (*) | 0.96 | (**) | 0.95 | (*) | 0.83 | (0.1) | 0.89 | (*) | 1.00 | (**) | |
| P | 0.93 | (**) | 0.98 | (**) | 0.93 | (**) | 0.93 | (**) | 0.90 | (**) | 0.96 | (**) | |
Tab. 1 shows the complete comparison across 2 semantic segmentation tasks, 8 target datasets and 2 microscopy modalities: mitochondria segmentation in 4 EM datasets and nuclei segmentation in 4 LM datasets. The table shows the mean and standard deviation of correlation scores across the set of task-specific target datasets (see Supp. for full tables).
We test CTE with a “hard” (CTE-NHD) and a “soft” (CTE-EI) consistency score, under input (Gauss) and feature space (DropOut) perturbations. In all cases, CTE demonstrates a strong correlation with segmentation performance. The other unsupervised method, Transfer Score (TS), underperforms across all datasets. Supervised metrics fail to take advantage of the target domain labels and do not reliably correlate with the ranking based on F1-score. As they were originally proposed to predict the final fine-tuned performance ranking, we also evaluate them on that task, observing the same lack of correlation as for zero-shot transfer ranking (see Supp.). The only exception is LEEP. It compares source-model predictions with target labels, building an empirical mapping between them and evaluating how well this mapping explains the target data. In direct-transfer settings with shared label sets, LEEP effectively reduces to the average log-likelihood of the correct target class, serving as a proxy for how well the source model’s predictive space separates target labels. Of note, LEEP is also the only baseline that—similar to CTE—operates in the predicted output space and does not make assumptions on the feature space geometry. Interestingly, LEEP is able to correlate for some target datasets (e.g., Hmito), but fails to correlate on average. The key difference between LEEP and LEEP is that the latter fits a Gaussian Mixture Model to the feature space in order to predict pseudo labels; however, it often struggles to converge for the segmentation models.

### v3

@@ 3.2 Our Consistency Model Ranking (CMR)
Considering a target dataset , we rank the transfer performance of a set of models on by measuring pixelwise prediction consistency between perturbed and unperturbed predictions (Fig.˜1). Firstly, we propose to extend the Effective Invariance (EI) [deng_strong_2022], originally formulated for classification tasks, to semantic segmentation. CMR-EI measures the per-class invariance of a single segmentation prediction to test-time perturbations,
@@
where and represent the pixelwise softmax and thresholded output of a model for class , with , the perturbed equivalents. To counteract high-class imbalance in biomedical data and allow for multiclass problems, we calculate a foreground restricted per-class consistency score, limiting iteration to the pixel set , thus preventing background pixels from dominating. CMR-EI is a “soft” metric as it utilises both output consistency and confidence, introducing a reliance on confidence calibration which is not guaranteed in transfer [ovadia_can_2019]. We therefore also propose a “hard” consistency measure based on the thresholded per-class binary model output. Using the normalised Hamming distance (NHD), we count the number of foreground restricted pixelwise label changes between the perturbed and unperturbed predictions,
This is equivalent to the , but considering pixel flips adds intuition. The consistency score for a single prediction is the average of the per-class scores.
For instance segmentation, direct individual pixel comparison is impossible as instance labels are arbitrary. Instead, we measure the consistency between and via the agreement of paired pixels in each prediction, as measured by the Rand Index [rand_objective_1971, unnikrishnan_measures_2005]. Intuitively, if a pair of pixels is assigned to a single instance in , then to be consistent, this should also be true in . To account for class imbalance, we use the foreground-restricted Adapted Rand Score (ARS) [arganda-carreras_crowdsourcing_2015],
where is the proportion of pixels that belong to instance in and instance in , restricted to the pairwise union of foreground pixels in and . represents the proportion of pixels with label in and is similarly defined for . Thus, is the proportion of pixel pairs belonging to a single instance in both and . From the perspective of either prediction, Eq.˜3 can then be considered as the weighted harmonic mean between two terms:
penalises split disagreements and is the proportion of pixels pairs in one instance in given they are in one instance in . penalises merge disagreements and is the proportion of pixel pairs in one instance in given they are in one instance in . The weighting is used by default. Following convention, every background pixel is treated as an instance containing a single pixel. The consistency score for a single transfer is then calculated as:
@@
where is calculated using one of the above metrics (CMR-EI, CMR-NHD, CMR-ARS), and we average over perturbations. The final ranking within is the descending order of .
@@ 3.3 Model Perturbation
Perturbations can be applied to inputs via TTA or directly to feature space, provided model access. We aim to assess model decision boundaries via target prediction consistency. This differs from evaluating the model’s learned invariance to perturbations within the source domain (e.g., due to training augmentations). To examine this, we rank models trained both with and without augmentation, showing perturbations still expose informative variations on target data. Perturbations must remain tolerable [mi_training-free_2022]: if too strong, even good models become inconsistent, while if overly weak all models stay stable. Empirically, we find a broad range of perturbations enable effective ranking via CMR (see Supp.).
Input Space Perturbations. We explored several popular image transformations: additive Gaussian noise, gamma correction, and changes in brightness and contrast. In our experiments, we found similar performance across all transformations. Thus, for brevity, in the main text we mostly only report Gaussian noise, with strength controlled by . See Supp. and Tab.˜2 for additional TTAs.
Feature Perturbation. Perturbations can be applied directly to intermediate feature layers, while keeping the network weights frozen. Motivated by findings in semi-supervised learning [ouali_semi-supervised_2020] and uncertainty estimation [mi_training-free_2022], we apply test-time dropout [tompson_efficient_2015, ledda_dropout_2023] (TTD). Specifically, we apply TTD (i) across all layers of the network, (ii) only at the bottleneck, where information is the most redundant [he_reshaping_2014] (Tabs.˜1 and 4) or (iii) at the bottleneck and skip connection layers (Tab.˜2).
@@ 4.1 Datasets
We evaluate our approach on a range of public segmentation datasets, covering six distinct tasks and many modalities. For each model ranking run, we fix one target dataset and evaluate all task-specific models (see Supp. for more details).
@@
Electron Microscopy (EM), Mitochondria. We use four semantic segmentation datasets with neural tissues from different species and EM modalities: [lucchi_learning_2013, franco-barranco_current_2023, phelps_reconstruction_2021]. We also reformulated these datasets as a per-patch binary classification task, where a positive label indicates a mitochondrion is present. This auxiliary task allows us to disentangle biomedical domain effects [chaves_performance_2023] from the intrinsic suitability of classification-based approaches for semantic segmentation.
Light Microscopy (LM), Nuclei. We use nine nuclei datasets for instance and semantic segmentation, taken from a variety of species and LM modalities: [vijayan_deep_2024, von_chamier_democratising_2021, kromp_annotated_2020, hawkins_rescu-nets_2024, ljosa_annotated_2012, caicedo_nucleus_2019, arvidsson_annotated_2023, de_helacytonuc_2024, leal-taixe_benchmark_2019] (see Supp. for detailed source target splits).
@@
Light Microscopy, Cells. We consider four instance cell segmentation datasets, taken from fruit fly, Arabidopsis thaliana and human samples: [leal-taixe_benchmark_2019, wolny_accurate_2020, willis_cell_2016, pape_microscopy-based_2021].
@@
Cone-Beam Computed Tomography (CBCT), ToothFairy2. We use the ToothFairy2 datasets taken from the MICCAI2024 challenge [2024TMI, 2025CVPR], comprising 3D CBCT volumes of the human jaw with 42 labelled classes. Model ranking is assessed on a held-out test set acquired at another institution.
4.2 Models and Correlation Evaluation Metrics
@@
Our diverse set of models include self-trained and pre-trained candidates. For classification: ResNet18/50 [he_deep_2015], DenseNet121 [huang_densely_2018], MobileNetV2/V3 [sandler_mobilenetv2_2018, Qian_MobileNetV3]. For binary semantic and instance segmentation: 2D U-Net [ronneberger_u-net_2015], Residual U-Net [lee_superhuman_2017], UNETR [hatamizadeh_unetr_2022], [archit_segment_2025], Cellpose-SAM [pachitariu_cellpose-sam_2025], and BMZ [ouyang_bioimage_2022] model, id: 5092850 ‘powerful-chipmunk’ [pape_microscopy-based_2021]. For multi-class semantic segmentation, we select top-models from the ToothFairy2 challenge [wang2025supervised] (Isensee et al. and Wang et al.) and further add architectures covering CNNs (2D nnU-Net, residual-encoder 3D nnU-Net [isensee2021nnu]), Transformers (SwinUNETR [hatamizadeh2021swin], TotalSegmentator [wasserthal_totalsegmentator_2023]) and State Space Models (UMamba [Ma2024], VMamba [vmamba]).
@@
For all ranking methods , we calculate the correlation between the metric ranking and the performance score ranking. As standard [agostinelli_how_2022], we evaluate the correlation scores using three complementary measures: Pearson correlation coefficient [benesty_pearson_2009], which captures linear relationships; Spearman’s rank correlation coefficient [spearman_proof_1904], which measures monotonicity; and Kendall tau [1972RankCM], which assesses ordinal association and is more sensitive to local pairwise rank disagreements. All three metrics range from with indicating perfect positive correlation. While a perfect should be monotonically related to , linear correlation is not required.
@@ 5.1 Semantic Segmentation
Previous work [agostinelli_how_2022, yang_pick_2023, pandy_transferability_2021] suggested that supervised transferability metrics developed for classification can be extended to semantic segmentation by sampling a class-balanced set of pixels and treating them as a classification dataset. However, for our biomedical datasets this approach fails: existing transfer metrics, even with supervision, do not correlate with either direct transfer performance (Tab.˜1(b)) or post–fine-tuning performance (see Supp.). To determine whether this limitation arises from the biomedical domain [chaves_performance_2023] or from the shift to segmentation, we additionally evaluated the metrics on mitochondria classification. Fig.˜2 shows model performance (F1) against each metric for a single target dataset (EPFL [lucchi_learning_2013]) for both classification and semantic segmentation. While the metrics correlate well with classification performance (violet crosses), they fail for segmentation (green circles), where only CMR and supervised LEEP remain reliable.
| Metric | Mito | Nuclei | |||||
| K | S | P | K | P | |||
| CMR-EI (Gauss) | 0.77 | 0.85 | 0.83 | 0.62 | 0.77 | 0.98 | |
| ±0.1 | ±0.1 | ±0.1 | ±0.1 | ±0.1 | ±0.0 | ||
| CMR-NHD (Gauss) | 0.71 | 0.83 | 0.79 | 0.69 | 0.82 | 0.97 | |
| ±0.1 | ±0.1 | ±0.1 | ±0.1 | ±0.1 | ±0.0 | ||
| CMR-EI (DropOut) | 0.73 | 0.86 | 0.86 | 0.74 | 0.85 | 0.90 | |
| ±0.1 | ±0.1 | ±0.1 | ±0.2 | ±0.1 | ±0.1 | ||
| CMR-NHD (DropOut) | 0.69 | 0.85 | 0.84 | 0.71 | 0.84 | 0.62 | |
| ±0.1 | ±0.1 | ±0.1 | ±0.2 | ±0.1 | ±0.4 | ||
| TS | 0.25 | 0.30 | 0.23 | 0.02 | 0.10 | -0.09 | |
| ±0.2 | ±0.3 | ±0.3 | ±0.3 | ±0.4 | ±0.4 | ||
| NuNo | 0.17 | 0.20 | 0.13 | 0.09 | 0.17 | 0.08 | |
| ±0.2 | ±0.3 | ±0.3 | ±0.5 | ±0.5 | ±0.5 | ||
| Dispersion | -0.03 | -0.07 | -0.17 | -0.18 | -0.29 | -0.19 | |
| ±0.0 | ±0.1 | ±0.1 | ±0.3 | ±0.4 | ±0.5 | ||
| Metric | Mito | Nuclei | |||||
| K | S | P | K | S | P | ||
| CCFV | -0.11 | -0.16 | -0.18 | -0.03 | -0.05 | -0.16 | |
| ±0.1 | ±0.1 | ±0.1 | ±0.3 | ±0.4 | ±0.5 | ||
| NLEEP | 0.41 | 0.49 | 0.39 | 0.11 | 0.12 | 0.32 | |
| ±0.4 | ±0.5 | ±0.3 | ±0.1 | ±0.1 | ±0.2 | ||
| LEEP | 0.89 | 0.96 | 0.97 | 0.67 | 0.77 | 0.94 | |
| ±0.1 | ±0.0 | ±0.0 | ±0.4 | ±0.3 | ±0.0 | ||
| GBC | 0.39 | 0.52 | 0.47 | -0.13 | -0.16 | -0.08 | |
| ±0.1 | ±0.1 | ±0.2 | ±0.3 | ±0.4 | ±0.4 | ||
| LogME | 0.08 | 0.10 | 0.06 | -0.20 | -0.28 | -0.49 | |
| ±0.2 | ±0.3 | ±0.3 | ±0.3 | ±0.4 | ±0.4 | ||
| NCTI | 0.21 | 0.28 | 0.37 | 0.29 | 0.25 | 0.30 | |
| ±0.2 | ±0.2 | ±0.2 | ±0.5 | ±0.6 | ±0.3 | ||
| RegHscore | 0.24 | 0.34 | 0.37 | 0.35 | 0.36 | 0.34 | |
| ±0.2 | ±0.4 | ±0.3 | ±0.4 | ±0.4 | ±0.3 | ||
Tab.˜1 reports the mean correlation across sets of 4 target datasets for mitochondria segmentation in EM [lucchi_learning_2013, franco-barranco_current_2023, phelps_reconstruction_2021] and nuclei segmentation in LM [von_chamier_democratising_2021, caicedo_nucleus_2019, ljosa_annotated_2012, arvidsson_annotated_2023] (see Supp. for per-dataset results). We evaluate CMR using both “hard” (CMR-NHD) and “soft” (CMR-EI) consistency metrics under input (Gauss) and feature (DropOut) perturbations. In all cases, CMR shows strong correlation with segmentation F1, outperforming the other unsupervised baselines—Transfer Score (TS), NuNo, and Dispersion. Most supervised transferability metrics also fail despite access to target labels, with LEEP as the only consistent exception. Most baseline methods operate in feature space and rely on assumptions about feature space geometry, such as inter-class separability or intra-class compactness [yang_can_2024, xie_importance_2023, deng_confidence_2023, pandy_transferability_2021, wang_how_2023]. The contrast between LEEP and LEEP, which adds feature-space reasoning, illustrates the difficulty of extending such assumptions from classification to segmentation. In contrast, CMR operates directly in the output space, comparing predictions rather than labels and thus enabling fully unsupervised ranking. Among the evaluated approaches, CMR provides the most reliable method for ranking models in practical repository settings.